OpenStudy (anonymous):

Written in slope-intercept form, the equation x + 2y = 4 is?

OpenStudy (zpupster):

y=mx+b subtract x from bothsides and divide both sides by 2

OpenStudy (mokeira):

y=mx+c rearrange 2y=-x+4 divide by 2 and get answer

OpenStudy (anonymous):

She will get confused with b and c

OpenStudy (anonymous):

Simply write your equation in this form ; y = .......

OpenStudy (anonymous):

slope is -1/2?

OpenStudy (zpupster):

yes

OpenStudy (zpupster):

y-intercept=??

OpenStudy (anonymous):

2

OpenStudy (zpupster):

good job!!

OpenStudy (anonymous):

thank you @zpupster :)

OpenStudy (zpupster):

you're welcome